### v2

# Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication

Abstract
By analyzing the spin-group symmetry of magnons, we establish two generalized Onsager’s relations in the magnon Hall effect, which reveals the rich and complicated structures of the magnon Berry curvature in the parameter space of different types of the exchange coupling. As an important consequence, we find that the isotropic exchange coupling is equally important as the Dzyaloshinskii-Moriya (DM) interaction in determining the magnon Hall effect, and it can transform to the latter under proper gauge transformation. Moreover, the diagonal and off-diagonal part of the symmetric anisotropic exchange coupling together can support the magnon Hall effect without the DM interaction. Both features are exemplified in ferromagnetic honeycomb lattices. Our work lays the ground for decoding the coupling between magnon transport and different types of exchange interactions.
Introduction.—As potential carriers in energy and spin transport, magnons, elementary excitations of spin orders, have drawn much attention in recent years. One outstanding example is the magnon Hall effect [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], which generates a thermal current perpendicular to the applied temperature gradient. The corresponding thermal conductivity can be related to the Berry curvature of magnon in the momentum space [5, 2, 1]. As a fundamental rule, the Onsager’s reciprocal relation[18, 19] restricts the structure of the transport coefficient. In the electronic counterpart, i.e., the anomalous Hall effect[20], it enforces the corresponding conductivity to be an odd function of the spin orientation, reveals the intimate relation between the band geometry and the order parameter, and hence serves as a guiding principle for experimental detection. In the magnonic case, however, besides separate case studies, a comprehensive understanding of the Onsager’s relation is still missing.
The difficulty lies in the identification of the critical parameter in the magnon transport. It is widely believed that the Dzyaloshinskii-Moriya (DM) interaction[21, 22] serves as an order parameter, associated with the breaking of the effective time-reversal symmetry consisting of a spin rotation and real time-reversal operation [4, 23, 17], so that the thermal Hall conductivity is an odd function of the DM interaction. However, it has been long overlooked that due to the spin-group symmetry, the choice of the spin rotation axis contains large gauge freedom, which essentially mixes the isotropic, anisotropic and antisymmetric part of the exchanging coupling. This complexity has been implicitly envisioned in recent studies which show that the anisotropic exchange coupling [24, 25] may also induce a nontrivial magnon Hall effect [7]. To accurately identify the critical parameter in the magnon Hall effect, a detailed analysis of the effective time reversal symmetry from the spin-group perspective is required.
In this work, we provide such a theory and derive two generalized Onsager’s relations. For any periodic magnetic textures, we use a local formulation of the magnonic Hamiltonian. We then identify the group $\tilde{G}$ of full effective time reversal operations (see Eq. (8)), inherited from the spin-only group[26, 27, 28, 29, 30, 31, 32, 33] in the original magnetic lattice. We find two types of effective time reversal operations, i.e., the uniform one and non-uniform one, from which two Onsager’s relations are established, revealing the rich and complicated dependence of the magnon Berry curvature on various exchange couplings. As important consequences, we find that the isotropic exchange coupling is as important as the DM interaction in determining the magnon Hall effect, and that the off-diagonal and diagonal part of the anisotropic exchange coupling can replace the role of the isotropic exchange coupling and DM interaction, respectively. Such findings are exemplified in magnons on ferromagnetic honeycomb lattices.
Local formulation of magnons.—We first set up the general framework for discussing the effective time reversal operation and its implication. Without loss of generality, we assume that the spin texture still possesses the lattice translation symmetry and that there can be arbitrary number of spins in a unit cell, each of which can point along different directions in equilibrium. An example of such system is shown in Fig.1. Instead of choosing a global coordinate system for the spin texture, we adopt the following rule: we define local coordinate systems on different lattice sites such that the $z$-axis of the local frame always points along the equilibrium direction of the local spin order, as shown in Fig.1. Due to the periodicity of the spin texture, the local coordinate frame defined in this way is also periodic.
We now write down the general spin Hamiltonian within this local frame. As the topic in concern is the transport of the spin wave, we focus on the Hamiltonian of the fluctuation around the equilibrium spin texture and ignore how such texture is achieved. For a stable spin texture, it is required that the energy is quadratic and positive with respect to any fluctuation. With this in mind, the spin Hamiltonian should take the following quadratic form:
$H=\frac{1}{2}\sum_{n,m}{\sum_{i,j}^{N}\sum_{ab}{\bm{S}_{ni}^{a}J^{ab}_{ni,mj}\bm{S}^{b}_{mj}}},$ (1)
where $S_{ni}^{a}$ is the $a$-th component of the local spin on the $i$-th lattice site in the $n$-th unit cell, and $J^{ab}_{ni,mj}$ contains all exchange interactions between local spins. Moreover, $J^{ab}_{ni,mj}=0$ when one of the $ab$ indices takes $z$ and the other one takes $x$ or $y$. For definiteness, we assume that there are $N$ different spins in a unit cell. It is clear that Eq. (1) contains all the quadratic forms of the fluctuation.
Figure 1: Configuration of a general magnetic system. Different colors of spins represent different sublattices and spins belong to the same sublattice have same orientation and local coordinate.
We now put the spin Hamiltonian into the canonical form. Through the Fourier transform and the Holstein-Primakoff transformation [34], the spin Hamiltonian can be expressed in the momentum space using the Nambu basis up to a constant: $H=\sum_{\bm{k}}\psi_{\bm{k}}^{\dagger}\mathcal{H}_{\bm{k}}\psi_{\bm{k}}$ with
$\mathcal{H}_{\bm{k}}=\frac{S}{2}\begin{bmatrix}\mathbf{A}(\bm{k})&\mathbf{B}(\bm{k})\\ \mathbf{B}^{*}(-\bm{k})&\mathbf{A}^{*}(-\bm{k})\end{bmatrix}.$ (2)
Here $\psi_{\bm{k}}=(a_{1,\bm{k}},\cdots,a_{N,\bm{k}},a^{\dagger}_{1,-\bm{k}},\cdots,a^{\dagger}_{N,-\bm{k}})^{T}$ represent the Nambu basis. The block of Eq. (2) can be expressed by $J^{ab}_{ni,mj}$ [35, 33, 14]
$A_{ij}(\bm{k})=$ $\sum_{m,n}{\frac{1}{2}\left(J_{ni,mj}^{+}-iD_{ni,mj}\right)}e^{i\bm{k}\cdot\left(\bm{R}_{mj}-\bm{R}_{ni}\right)}$
$-\delta_{ij}\sum_{mn\ell}{J_{ni,m\ell}^{zz}},$ (3)
$B_{ij}(\bm{k})=$ $\sum_{m,n}{\frac{1}{2}\left(J_{ni,mj}^{-}+i\Gamma_{ni,mj}\right)}e^{i\bm{k}\cdot\left(\bm{R}_{mj}-\bm{R}_{ni}\right)},$ (4)
where $\bm{R}_{ni}$ is the position of the $i$-th spin in the $n$-th unit cell, $J^{+}_{ni,mj}=J_{ni,mj}^{xx}+J_{ni,mj}^{yy}$, $J^{-}_{ni,mj}=J_{ni,mj}^{xx}-J_{ni,mj}^{yy}$, $D_{ni,mj}=J_{ni,mj}^{xy}-J_{ni,mj}^{yx}$, and $\Gamma_{ni,mj}=J_{ni,mj}^{xy}+J_{ni,mj}^{yx}$. By summing over the unit cell index, we can further define the following symbols according to Eq. (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication): $A_{ij}=J_{ij}^{+}-iD_{ij}-\delta_{ij}\sum_{mn\ell}J_{ni,m\ell}^{zz}$ and $B_{ij}=J_{ij}^{-}+i\Gamma_{ij}$. One can further check that $A(\bm{k})=A^{\dagger}(\bm{k})$ and $B(\bm{k})=B^{T}(-\bm{k})$. As a result, $\mathcal{H}$ is Hermitian.
Interestingly, the definition of the intermediate quantities $J^{\pm}$, $D$ and $\Gamma$ agrees with the decomposition of the exchange coefficient. With fixed lower indices, $J^{ab}_{ni,mj}$ is a rank-2 tensor, which can then be decomposed into three parts in the $xy$ subspace: the scalar part, also referred to as the isotropic exchange interaction, is represented by $J^{+}$, the antisymmetric part, also referred to as the DM interaction, is represented by $D$, the traceless symmetric part is represented by $J^{-}$ and $\Gamma$. We shall emphasize the $\pi/2$ phase difference between $J^{+}$ and $D$, and $J^{-}$ and $\Gamma$, and the fact that $J^{+}$ and $D$ work in the particle or hole sector but $J^{-}$ and $\Gamma$ mix particles with holes in the magnon Hamiltonian.
The Hamiltonian in Eq. (2) can be diagonalized by using para-unitary transformation[36, 37, 38, 39]
$T_{\bm{k}}^{\dagger}\mathcal{H}_{\bm{k}}T_{\bm{k}}=\begin{bmatrix}E_{\bm{k}}&\\ &E_{\bm{-k}}\end{bmatrix},$ (5)
where $E_{\bm{k}}=\mathrm{diag}(\varepsilon_{1,\bm{k}},\varepsilon_{2,\bm{k}},\cdots,\varepsilon_{N,\bm{k}})$. The structure of the eigen energy is due to particle-hole symmetry. In order to maintain bosonic commutation rules, the transformation matrix $T_{\bm{k}}$ has to satisfy a special orthonormal condition $T_{\bm{k}}^{\dagger}\Sigma_{z}T_{\bm{k}}=\Sigma_{z}$, where $\Sigma_{z}=\sigma_{z}\otimes\mathbb{I}_{N}$ with $\sigma_{z}$ being the Pauli matrix and $\mathbb{I}_{N}$ being the $N\times N$ unit matrix.
Within this framework, the thermal Hall conductivity of non-interacting magnons is given by [2, 5]
$\kappa_{\mu\nu}=-\frac{k_{B}^{2}T}{\hbar(2\pi)^{2}}\sum_{n=1}^{N}{\int_{BZ}{c_{2}\left[\rho\left(\varepsilon_{n,\bm{k}}\right)\right](\Omega_{\mu\nu})_{n,\bm{k}}\mathrm{d}^{2}k}},$ (6)
where $c_{2}(\rho)=(1+\rho)(\mathrm{log}\frac{1+\rho}{\rho})^{2}-(\mathrm{log}\rho)^{2}-2\mathrm{Li}_{2}(-\rho)$ with $\mathrm{Li}_{2}(z)$ being the polylogarithm function, $\rho_{n}$ is Bose distribution function of the $n$-th band, and $\Omega_{n,\bm{k}}$ is Berry curvature [38]:
$(\Omega_{\mu\nu})_{n,\bm{k}}=i\left[\Sigma_{z}\frac{\partial T_{\bm{k}}^{\dagger}}{\partial k_{\mu}}\Sigma_{z}\frac{\partial T_{\bm{k}}}{\partial k_{\nu}}\right]_{nn}.$ (7)
Generalized Onsager’s Relation.—At the heart of the Onsager’s relation is the time reversal operation $T$. In the electronic Hall-type conductivity, on one hand, $T$ connects systems with opposite magnetic order; on the other hand, $T$ flips the sign of the Hall conductivity. The Onsager’s relation then states that the Hall conductivity is an odd function of the magnetic order[20, 40].
In the magnonic degree of freedom, however, the situation is quite different. Neither the Hamiltonian of the fluctuation in Eq. (1) nor the Berry curvature in Eq. (7) can be flipped by the time reversal operation $T$. Therefore, using $T$ alone cannot offer any useful information about the structure of the magnon Hall conductivity.
A practical generalization is the effective time reversal operation $\mathcal{T}=TC_{2x}^{s}$ [41, 42] where the rotation only acts on spin and flips the spin order. Since $\mathcal{T}$ can flip both the magnon Berry curvature and the DM interaction, it is usually assumed that the DM interaction plays the role of ‘magnetic order’ and is indispensable in the magnon Hall effect[3, 12, 4, 9, 10, 11, 13, 14, 16, 8, 15]. We shall emphasize that in previous study[23, 4, 17], this analysis is applied in ferromagnets and coplanar antiferromagnets where a global $C_{2}$ spin rotation axis exists.
In the general setting, a significant issue arises: what is the precise meaning of the $C_{2}$ spin rotation axis? For non-coplanar spin textures, there is no global spin rotation axis that can simultaneously flip all the spin orders. The best one can define is the local spin rotation axis that differ site by site, as employed in our framework and illustrated in Fig. 1.
The locally defined spin rotation axis makes prominent an elemental issue, i.e., the gauge freedom in choosing the local spin-x direction. At each site in a unit cell, only the spin-z direction is well defined, and any direction perpendicular to the $z$-th direction can work equally well as the $x$ direction. The spin-x axis on different sites can have no connection. This gauge freedom should also emerge in ferromagnetic and coplanar antiferromagnetic crystals.
Mathematically, the above issue is related to the spin group of the magnetic crystals. In ferromagnets, the magnetic order is subject to the spin-only group $G=\infty 2^{\prime}$[26, 27, 28, 29, 30, 31, 32, 33], where $C_{2}$ rotation axis is perpendicular to the $C_{\infty}$ axis (i.e., the direction of the ferromagnetic order). For general spin structures, the spin group obeyed by the magnetic order should at least contain the direct product of the subgroup for each one of the spin in a unit cell
$\tilde{G}=\bigotimes_{i=1}^{N}G_{i},$ (8)
where $G_{i}=\infty 2^{\prime}$ with the axis defined in the local spin frame. We emphasize that more symmetry operations can appear that connect the spin orders on different sites. The group $G$ is just the common subgroup obeyed by any periodic spin textures. Although $\tilde{G}$ has been derived previously[32, 33], its implication in magnon transport is still unclear.
This group structure greatly restricts the magnon Hall effect. To see this, we make a gauge transformation that rotates the spin axis on the i-th site about the local $z$-axis by an angle $\phi_{i}$ with i going through all the local spin orders in a unit cell. As a result, we find that $A_{ij}\rightarrow A_{ij}e^{i(\phi_{i}-\phi_{j})}$ and $B_{ij}\rightarrow B_{ij}e^{i(\phi_{i}+\phi_{j})}$. Since the magnon Hall conductivity should not depend on any specific choice of $\phi_{i}$, the model parameters $J_{ni,mj}^{\pm}$, $D_{ni,mj}$ and $\Gamma_{ni,mj}$ should only appear in the form of gauge-independent clusters, the number of which is at least $N$-less than the total number of parameters in Eq. (2).
Since our goal is to study the Onsager’s relation, we shall focus on the effective time reversal operation in this spin group, i.e., $\mathcal{T}_{x}=TC_{2x}^{s}$ and $\mathcal{T}_{y}=TC_{2y}^{s}$. Although physically equivalent, the arbitrary substitution of one with the other yields distinct constraints on the model parameters.
We start with choosing $\mathcal{T}_{x}$ identically on each spin in a unit cell, which is the uniform effective time reversal operation. We emphasize that although the mathematical symbol is the same, $\mathcal{T}_{x}$ on different spins can involve $C_{2x}$ rotations along different axis in the lattice frame. In the magnon case, the effective time reversal operation $\mathcal{T}_{x}$ is the spinless-version of the time reversal operation, and can be identified as $\mathcal{T}_{x}=-K$ where $K$ stands for complex conjugation. We then find that[35] $\mathcal{T}_{x}\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})\mathcal{T}_{x}^{-1}=\mathcal{H}_{-\bm{k}}^{*}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})=\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$. The last equality holds by directly evaluating the Hamiltonian at $-\bm{k}$. By calculating the Berry curvature in the original parameter sets $(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})$ but with the time reversal operator and in the new parameter sets $(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$, respectively, we find that
$(\Omega_{\mu\nu})_{n,\bm{k}}(J_{ij}^{\pm},D_{ij},\Gamma_{ij})=-(\Omega_{\mu\nu})_{n,-\bm{k}}(J_{ij}^{\pm},-D_{ij},-\Gamma_{ij}),$ (9)
We then obtain the first Onsager’s relation[35]:
$\kappa_{\mu\nu}(J_{ij}^{\pm},-D_{ij},-\Gamma_{ij})=-\kappa_{\mu\nu}(J_{ij}^{\pm},D_{ij},\Gamma_{ij}).$ (10)
The effective time reversal symmetry can also be chosen in a non-uniform manner. We classify the spins in a unit cell into two groups, such that the time reversal operation on spins in the first group is $\mathcal{T}_{x}$ while that on spins in the remaining group is $\mathcal{T}_{y}$. There are consequently two types of parameters $J_{ij}^{\pm}$, $D_{ij}$, and $\Gamma_{ij}$, i.e., those with $i$ and $j$ within the same group, labeled by $J_{ij}^{\pm,intra}$, $D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}$ and those with $i$ and $j$ belonging to different groups, labeled by $J_{ij}^{\pm,inter}$, $D_{ij}^{inter}$, and $\Gamma_{ij}^{inter}$. Since $\mathcal{T}_{y}=K$, we then find that the total effective time reversal symmetry reads as $\mathcal{T}=(-KI_{n_{1}})\oplus(KI_{n_{2}})$ with $n_{1}$ and $n_{2}$ being the number of spins in the first and second group, respectively. The effect of $\mathcal{T}$ on the intra-group parameters is the same as in the uniform case, i.e., $J_{ij}^{\pm,intra}\rightarrow J_{ij}^{\pm,intra}$, $D_{ij}^{intra}\rightarrow-D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}\rightarrow-\Gamma_{ij}^{intra}$. Due to the relative minus sign between $\mathcal{T}_{x}$ and $\mathcal{T}_{y}$ in $\mathcal{T}$, the effect on the inter-group parameters is opposite: $J_{ij}^{\pm,intra}\rightarrow-J_{ij}^{\pm,intra}$, $D_{ij}^{intra}\rightarrow D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}\rightarrow\Gamma_{ij}^{intra}$. We can then obtain the second Onsager’s relation[35]:
$\kappa_{\mu\nu}(J_{ij}^{\pm,intra},-D_{ij}^{intra},-\Gamma_{ij}^{intra},-J_{ij}^{\pm,inter},D_{ij}^{inter},\Gamma_{ij}^{inter})$
$=-\kappa_{\mu\nu}(J_{ij}^{\pm,intra},D_{ij}^{intra},\Gamma_{ij}^{intra},J_{ij}^{\pm,inter},D_{ij}^{inter},\Gamma_{ij}^{inter})$ (11)
The two Onsager’s relations in Eq. (10) and (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication) are main results of this work. They reveal complicated dependence of the magnon Berry curvature on different types of exchange couplings. The first Onsager’s relation can obviously reproduce the usual understanding that the magnon Hall effect is induced by the DM interaction with vanishing anisotropic parameter $\Gamma_{ij}$. Besides that, the magnon Hall effect can have much richer structures.
Implications.—One direct consequence of the second Onsager’s relation is that the isotropic exchange interaction can be as important as the DM interaction in determining the magnon Hall effect. In fact, the role of the isotropic exchange and that of the DM interaction are even interchangeable by a proper gauge transformation. On one hand, with vanishing anisotropic exchange coupling, the two Onsager’s relations suggest that the magnon Hall conductivity is an odd function of both the DM interaction and the isotropic exchange coupling. On the other hand, from Eq. (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication), an additional $\pi/2$ phase factor in $A_{ij}$ inverts its real and imaginary part and hence leads to $J_{ni,mj}^{+}\rightarrow-D_{ni,mj}$ and $D_{ni,mj}\rightarrow J_{ni,mj}^{+}$. Physically, such a phase factor amounts to rotating the spin frame at either one of the $i$ and $j$ sites around the local $z$ axis by $\pi/2$, which should not generate any measurable difference.
Figure 2: (a) Honeycomb lattice with in-plane magnetization pointing $45^{\circ}$ from the horizontal line. (b) The thermal conductivity as a function of $J_{N}$ and $D$. In (b), $J_{A}=0.1\sqrt{J_{N}^{2}+D^{2}}$, $J_{B}=0.15\sqrt{J_{N}^{2}+D^{2}}$; the blue and red curve have positive and negative sign, respectively.
As a concrete example, we consider a ferromagnetic honeycomb lattice with nearest and next nearest exchange coupling, DM interaction and easy-plane anisotropy. The corresponding magnetic Hamiltonian reads as
$H=$ $-\sum_{\langle i,j\rangle}{J_{N}\bm{S}_{i}\cdot\bm{S}_{j}}-\sum_{\langle\langle i,j\rangle\rangle}{\sum_{\mu=A,B}{J_{\mu}\bm{S}_{i}^{\mu}\cdot\bm{S}_{j}^{\mu}}}$ (12)
$+\sum_{\langle i,j\rangle}{D\left(\bm{R}_{ij}\times\hat{\bm{z}}\right)\cdot\left(\bm{S}_{i}\times\bm{S}_{j}\right)}.$
where $\bm{R}_{ij}$ is the unit vector from site $i$ to $j$. In the nearest neighbor exchange coupling, when $J_{A}=J_{B}$, the chiral symmetry $\mathcal{C}=\sigma_{x}K$ is respected, and the system is hence gapless with vanishing magnon Hall effect. Moreover, an external magnetic field is implied which stabilizes the in-plane ferromagnetic order, as shown in Fig. 2(a).
In Fig. 2(b), we plot the Hall conductivity as a function of $J_{N}$ and $D$, with $J_{N}^{2}+D^{2}$ fixed. We find that $\kappa_{xy}$ has four lobes with alternating signs, resembling the angular dependence of $d_{xy}$ wave. As a result, $\kappa_{xy}$ is indeed an odd function of $J_{N}$ and $D$, consistent with the two Onsager’s relations. Such sign-change of the magnon Hall conductivity can potentially be observed in materials with ferromagnetic-antiferromagnetic transition.
The $d$-wave pattern further suggests the equivalence between $J_{N}$ and $D$. To see this, we note that each A site has three nearest-neighbor bonds, which we label by $1$, $2$, $3$, as shown in Fig. 2. The exchanging coupling parameter on each bond forms a two-dimensional vector, with $(J_{1},D_{1})=(J_{N},-D/\sqrt{2})$, $(J_{2},D_{2})=(J_{N},D\cos(\pi/12))$, and $(J_{3},D_{3})=(J_{N},-D\sin(\pi/12))$. A rotation of the local spin frame at site A then interchanges the role of $J_{i}$ and $D_{i}$: $(J_{i},D_{i})\rightarrow(D_{i},-J_{i})$ with $i=1,2,3$. More specifically, since $\kappa_{xy}$ should be unchanged under such rotation of local spin frame, only the relative orientation angles between the three $(J_{i},D_{i})$ survive. The leading order contribution to $\kappa_{xy}$ should then has the following pattern: $\kappa_{xy}=A_{1}(J_{1},D_{1})\times(J_{2},D_{2})+A_{2}(J_{1},D_{1})\times(J_{3},D_{3})+A_{3}(J_{2},D_{2})\times(J_{3},D_{3})$, where $A_{i}$ is a function of the magnitude of the three vectors $(J_{\ell},D_{\ell})$, and the cross product of two two-dimensional vector generates a pseudoscalar. Therefore, $\kappa_{xy}$ is a function of $J_{N}\cdot D$, consistent with the d-wave pattern and clearly demonstrates the equivalence between $J_{N}$ and $D$. Moreover, since $|D_{i}|<D$, $J_{i}^{2}+D_{i}^{2}$ is always slimmer in the $D$-direction, as can be observed in Fig. 2(b).
Another consequence of the two Onsager’s relations is that the DM interaction is not necessary for the magnon Hall effect. It can be fully replaced by the anisotropic exchange coupling. By setting $D=0$, Eq. (10) suggests that $\kappa$ is an odd function of the anisotropic exchange parameter $\Gamma$. Moreover, similar to the equivalence between $D$ and $J^{+}$, for the anisotropic exchange coupling, the off-diagonal part $\Gamma_{ni,mj}$ is equivalent to $J_{ni,mj}^{-}$, due to a rotation of the local spin frame at either $i$ or $j$ by $\pi/2$.
Figure 3: (a) The thermal conductivity as a function of $\Gamma$ and $J^{-}$. Parameters are $J^{+}=\pm 0.5J_{0}$ which changes sign as $\Gamma$ varies, $\sqrt{(J^{-})^{2}+\Gamma^{2}}=0.1J_{0}$, $J_{A}=0.05J_{0}$, and $J_{B}=0.075J_{0}$. The blue and red curve are positive and negative, respectively. (b) Symmetry requirements for DM interaction(top) and anisotropic exchange coupling(bottom).
To show this, we consider the following ferromagnetic honeycomb model with the magnetic order being along $z$-axis:
$H=\sum_{\langle i,j\rangle}{\bm{S}_{i}^{T}\mathcal{J}_{ij}\bm{S}_{j}}-\sum_{\langle\langle i,j\rangle\rangle}{\sum_{\mu=A,B}{J_{\mu}\bm{S}_{i}^{\mu}\cdot\bm{S}_{j}^{\mu}}}\,,$ (13)
where
$\mathcal{J}_{ij}=\left(\begin{matrix}-J^{+}+J^{-}&w_{ij}\Gamma&0\\ w_{ij}\Gamma&-J^{+}-J^{-}&0\\ 0&0&-J_{0}\\ \end{matrix}\right).$ (14)
where $\Gamma$, $J^{+}$ and $J^{-}$ are constants representing off-diagonal anisotropic, isotropic, diagonal anisotropic exchange couplings, respectively. $w_{ij}$ are weighing factors for the three nearest-neighbor bonds, and they are chosen to break the in-plane mirror symmetry. The next nearest neighbor term is added to break the chiral symmetry as before.
In Fig. 3 we plot the thermal Hall conductivity as a function of $\Gamma$ and $J^{-}$. One can find a similar $d$-wave pattern, showing that $\kappa_{xy}$ is an odd function of both $\Gamma$ and $J^{-}$, and the equivalence between $\Gamma$ and $J^{-}$. We shall emphasize that the anisotropic exchange coupling only affects $J^{-}$ and $\Gamma$ in our case, while the one in previous study[7] also serves as antisymmetric exchange interaction in the magnon Hamiltonian[35]. Therefore, our work unambiguously demonstrates that the anisotropic exchange alone can support the magnon Hall effect.
The anisotropic exchange coupling has completely different symmetry requirements from the DM interaction. The latter requires the breaking of inversion symmetry. In comparison, the former requires the breaking of mirror-x and mirror-y symmetry; the inversion symmetry can be retained. This then extends the material candidate for the magnon Hall effect.
Acknowledgements.
We acknowledge useful discussions with Shiyue Deng, Shu Li and Qingtao Zhao. This work is supported by the National Natural Science Foundation of China (12234017). Y. G. is also supported by the Innovation Program for Quantum Science and Technology (2021ZD0302802). The supercomputing service of USTC is gratefully acknowledged.